(1 point) The heights of women aged 20 to 29 follow approximately the N(64, 2.68) distribution. Men the same age have heights distributed as N(69.3 2.69). What percent of young women are taller than the mean height of young men?(1 point) Changing the mean and standard deviation of a Normal distribution by a moderate amount can greatly change the percent of observation in the tails. Suppose that a college is looking for applicants with SAT math scores 750 and above (a) In 2007, the scores of men on the math SAT followed the N(533, 116) distribution. What percent of men scored 750 or better? (b) Womens SAT math scores that year had the N(499, 110) distribution. What percent of women scored 750 or better?
Answer #1

Please post 1 question per post as per forum rules.

1.

Normal distribution has been given whose params are :

women:

mean = 64,
Stdev = 2.68

men:

mean =69.3
stdev = 2.68

%age of young women are taller than mean height of young men
= P(X>69.3)
= P(Z> (69.3-64)/2.68) = P(Z>1.978) = 1-0.976 = .024 or 2.4%

So, 2.69% of young women are taller than the mean height of young men